the sum of the squares of two consecutive even integers is 52. what are the integers? I believe the answers is 4^2 and 6^2 but I can't figure out how to write an equation to show how I figured it out, Please help!!!: the sum of the squares of two consecutive even integers is 52. what are the integers? I believe the answers is 4^2 and 6^2 but I can't figure out how to write an equation to show how I figured it out, Please help!!!
Answer by edjones(2169)

Let the smaller be x, the larger x+2
x^2+(x+2)^2=52
x^2+x^2+4x+4=52
2x^2+4x-48=0
2(x^2+2x-24)=0
(x+6)(x-4)=0
x=4
x+2=6
.
Ed
Question 148437: the sum of the squares of two consecutive even integers is 52. what are the integers? I believe the answers is 4^2 and 6^2 but I can't figure out how to write an equation to show how I figured it out, Please help!!!: the sum of the squares of two consecutive even integers is 52. what are the integers? I believe the answers is 4^2 and 6^2 but I can't figure out how to write an equation to show how I figured it out, Please help!!!
Answer by Earlsdon(3498)

Try this!
Let x be the first even integer, then (x+2) is the next consecutive even integer.
The sum of their squares is expressed by:
x^2+(x+2)^2 and this is equal to 52, so you can write the equation:
x^2+(x+2)^2 = 52 Simply this.
x^2+(x^2+4x+4) = 52 Combine like-terms and subtract 52 from both sides.
2x^2+4x-48 = 0 Now you have a quadratic equation that can be solved by factoring after dividing through by 2 to simplify it a bit.
x^2+2x-24 = 0 Factor.
(x-4)(x+6) = 0 from which you get:
x = 4 or x = -6
The integers are 4 and 6 or -6 and -4
Check:
4^2+6^2= 16+36=52 or...
(-6)^2+(-4)^2 = 36+16=52